OpenStudy (anonymous):

How can x + 8 = -x + 7 be set up as a system of equations?

jimthompson5910 (jim_thompson5910):

set each side equal to y y = x+8 y = -x+7

jimthompson5910 (jim_thompson5910):

to get back to x + 8 = -x + 7, you would... y = -x+7 start with the second equation x+8 = -x+7 ... replace y with x+8 (works because y = x+8)